Candy Chemist was asked to determine the concentration (M & X) of a new alcoholic beverage from the mass percent ethanol in aqueous solution, which was determined to be 15.0% (m/m), and the density, which was found to be0.984 g/mL. What is the molarity and mole fraction of ethanol (MM = 46.068 g/mol) in the beverage?
Answer #1

Molarity = 10*(mass%)*d/mol.wt
molarity = 10*15*0.984/46.068 = 3.204M

mole fraction

χ1 = (M x mw2) / [(M x mw2) + (d - M x mw1)]

where
mw1 = molecular weight solute
mw2 = molecular weight solvent
d = density solution
M = molarity
χ1 = mole fraction solute